# PAC-Bayes Analysis for Recalibration in Classification

### v1

@@ 4.2 Proposed recalibration algorithm based on PAC-Bayes bounds
[…]
However, minimizing this function is difficult since the ECE is neither smooth nor differentiable. To avoid this issue, we focus on the following fact: $\mathrm{ECE}(\eta_{v}(f_{w}),S_{\mathrm{re}})\leq\frac{1}{n_{\mathrm{re}}}\sum_{(X,Y)\in S_{\mathrm{re}}}\|\hat{Y}-\eta_{v}(f_{w}(X))\|_{2}^{2}$ in the multiclass classification  [^2] A similar relation holds for binary classification. The complete proof can be found in Appendix B.3.. This upper bound the Brier score [9], which is a continuous and differentiable calibration metric. Therefore, under the posterior $\tilde{\rho}(v)=\tilde{\rho}(v;\theta)$ parametrized by $\theta$, we can minimize the following PAC-Bayes-based objective by gradient-based optimization:
$\min_{\theta}\mathbb{E}_{\tilde{\rho}(v;\theta)}\frac{1}{n_{\mathrm{re}}}\sum_{(X,Y)\in S_{\mathrm{re}}}\|\hat{Y}-\eta_{V}(f_{w}(X))\|^{2}_{2}+\alpha\frac{\mathrm{KL}(\tilde{\rho}(v;\theta)\|\tilde{\pi})}{n_{\mathrm{re}}}.$ (15)
@@
$\min_{\theta}\mathbb{E}_{\tilde{\rho}(v;\theta)}\frac{1}{n_{\mathrm{re}}}\sum_{(X,Y)\in S_{\mathrm{re}}}(l_{\mathrm{acc}}(Y,\eta_{V}(f_{w}(X)))+\|\hat{Y}-\eta_{V}(f_{w}(X))\|_{2}^{2})+\alpha\frac{\mathrm{KL}(\tilde{\rho}(v;\theta)\|\tilde{\pi})}{n_{\mathrm{re}}}.$ (16)
Eq. (16) is also guaranteed through PAC-Bayes bounds in terms of the test accuracy and TCE (or CE) simultaneously because it can be derived by utilizing the union bound for Eq. (1) and Corollary 2. The detailed derivation can be seen in Corollary 5 in Appendix B.3.
We refer to the recalibration achieved by minimizing Eqs. (15) and (16) as PAC-Bayes recalibration (PBR) and summarize these overall procedures in Algorithm 1. As an example of a recalibration model in PBR, we can consider using the Gaussian process (GP) [25]. In this case, $\tilde{\rho}(v;\theta)$ is set as the multivariate Gaussian distribution, $\mathcal{N}(v;\mu,\Sigma)$, where $\mu$ and $\Sigma$ are the mean and covariance matrix, respectively. Furthermore, as discussed in Section 4.1, we can construct the GP prior $\tilde{\pi}$ using the output values of $f_{w}(x)$ for $x\in S_{\mathrm{re}}$ during training, where $w$ is learned using $S_{\mathrm{tr}}$. This is because such $\tilde{\pi}$ can be dependent on $S_{\mathrm{tr}}$ since $S_{\mathrm{tr}}$ and $S_{\mathrm{re}}$ are independent. In this way, we can construct a data-dependent prior, which is expected to yield a small KL value.
A similar GP-based recalibration model was also used by Wenger et al. [32]. In this study, we propose recalibration via variational inference (VI), i.e., minimizing the following objective function derived from the evidence lower bound (ELBO): $\mathbb{E}_{\tilde{\rho}(v;\theta)}\frac{1}{n_{\mathrm{re}}}\sum_{(X,Y)\in S_{\mathrm{re}}}l_{\mathrm{acc}}(Y,\eta_{V}(f_{w}(X)))+\mathrm{KL}(\tilde{\rho}(v;\theta)|\tilde{\pi})/n_{\mathrm{re}}$, where $l_{\mathrm{acc}}$ is defined as the softmax loss. Eq. (16) corresponds to this ELBO objective when $\alpha=1$ and the Brier score is removed. This fact indicates that PBR extends the GP-based method to have flexible KL regularization and Brier score loss to control calibration performance on the basis of the PAC-Bayes notion. In Section 6.2, we compare PBR and GP-based recalibration using the same recalibration model settings.
@@
6:  (Option) Calculate the test ECE: $\mathrm{ECE}(\eta_{V}(f_{w}),S_{\mathrm{te}})$ by setting $B=\lfloor n_{\mathrm{re}}^{1/3}\rfloor$.
7:  RETURN: $\eta_{V}(f_{w})$, ($\mathrm{ECE}(\eta_{V}(f_{w}),S_{\mathrm{te}})$)
5 Related work
In classification, some exciting studies have utilized the PAC-Bayes theory to analyze various metrics other than accuracy. Ridgway et al. [26] studied the AUC metric. Morvant et al. [22] analyzed the confusion matrix in multiclass classification. We extend this line of inquiry to the classification metric by focusing on calibration. Additionally, in terms of uncertainty quantification, Sharma et al. [28] analyzed the conformal prediction, which is a different metric for uncertainty quantification.
Gupta and Ramdas [11] and Kumar et al. [18] examined the ECE with binning in binary classification focusing on uniform mass binning, which divides intervals of the same mass. They also analyzed a recalibration method based on this binning method. In contrast, our analysis of the ECE in a multiclass setting, which has not yet been performed, focuses on the typical equal-width binning method. Sun et al. [29] also analyzed the binning-based recalibration for binary classification under a similar Lipschitz assumption. They have shown the same estimation bias order as in our results only when the histogram binning recalibration is performed, suggesting $\mathcal{O}(1/n^{1/3})$. This result is expected from the viewpoint of nonparametric regression, as discussed in Section 3.1.
The analysis of multiclass calibration is limited compared with the binary classification. Zhang et al. [38] investigated the kernel-based $CE_{K}$. However, their analysis does not address generalization or recalibration, which fundamentally distinguishes our approach. Moreover, Zhang et al. [38] remarked that the binning-based $CE_{K}$ suffers from the curse of dimensionality without a formal statement, whereas our Theorem 4 validates this phenomenon. Gruber and Buettner [9] discussed various calibration metrics and clarified that the TCE is a weaker calibration metric than to $CE_{K}$. On the other hand, our Theorem 4 enables the non-asymptotic bias analysis, showing that the naive estimator of $CE_{K}$ suffers from the curse of dimensionality compared with the TCE.
6 Experiments
In this section, we confirm the correlation between our bounds in Corollary 1 and the generalization performance of recalibration (Section 6.1). In Section 6.2, we discuss the effectiveness of our PBR in Section 4.2. Owing to page limitation, we report here the results of the multiclass classification experiments on the MNIST [19] and CIFAR-100 [15] datasets. All experimental results can be found in Appendix D. In addition, the details of our experimental settings, including model information, are summarized in Appendix C. We also set $B=\lfloor n_{\mathrm{re}}^{1/3}\rfloor$ for ECE evaluation on the basis of our findings.
6.1 Verification of our bounds
In this section, we empirically confirm the correlation between the KL divergence term in Eq. (13) and the generalization gap in the ECE. To achieve this, we conducted recalibration experiments using PBR presented in Eq. (15) with various constraints ($\alpha$) for the KL term. Then, we measured each KL value and ECE gap after recalibration by performing $10$-fold cross-validation, where we set $n_{\mathrm{re}}=1000$.
A portion of the results is shown in Figure 1. From these results, we can see that the KL and generalization gap of the ECE are somewhat correlated in terms of Pearson’s and Kendall’s rank-correlation coefficients, which validates our bound. It is also evident that the ECE gap is noisy. This is because the estimation bias of the ECE through binning has a very slow convergence rate of $\mathcal{O}(1/n_{\mathrm{re}}^{1/3})$, even under our optimal $B$. When using $n_{\mathrm{re}}=1000$, the noise due to the estimation bias can be $\mathcal{O}(1/10)$. This implies that the degree of correlation is likely to be buried in this noise if models achieve a small ECE of less than $0.1$. A simple way to reduce the noise is to increase the number of recalibration data points; however, this is impractical according to $\mathcal{O}(1/n_{\mathrm{re}}^{1/3})$ convergence rate.
6.2 Empirical comparison of recalibration methods
In this section, we compare our proposed PBR and existing standard or recent recalibration methods, temperature scaling [10] and GP calibration [32]. The optimizable parameter $\alpha$ in PBR is selected from $0\}$ , 0.01 , 0.1 , 0.25 , 0.5 , 0.75 , 0.9 , 1.0 } by grid search.
We summarize the results in Table 1. These results show that our method effectively improves calibration performance over existing methods, especially the GP-based recalibration method. We also found that, contrary to the report by Wenger et al. [32], temperature scaling performs better than GP-based recalibration in some cases in terms of the ECE with the optimal $B$. The failure to reproduce this result could be because Wenger et al. [32] reported the ECE with $B=100$. According to our analysis, this setting achieves the smallest estimation bias when $n_{\mathrm{te}}=\mathcal{O}(100^{3})$. However, since the $n_{\mathrm{te}}$ actually used is far smaller than this (see Table 2 in Appendix C), severe bias may occur.
Despite these instabilities, PBR appears to consistently improve the existing GP methods, at least in the case of the ECE under the optimal $B$ studied in this paper. The results of other experiments shown in Appendix D confirmed that GP-based methods, including PBR, do not consistently outperform standard recalibration methods such as temperature scaling. This indicates the need to select a recalibration model appropriate for the dataset and model used.
@@ 6.2 Empirical comparison of recalibration methods
Figure 1: KL vs ECE gap on MNIST (Multiclass; $1$-layer NN) and CIFAR-100 (Multiclass; ResNeXt-$29$ [34] and DenseNet-BC-190 [13]) dataset per various regularize parameters ($\alpha$).
@@
Table 1: Results of multiclass classification experiments on MNIST and CIFAR-100. The symbols ($\downarrow$) and ($\uparrow$) refer to lower and higher values indicating better performance, respectively.
Data Model Uncalibrate GP [32] PBR (Ours; Eq. (15)) PBR (Ours; Eq. (16))) Temp. Scaling
@@
MNIST XGBoost .$0036\pm.0003$$9785\pm.0005$$0038\pm.0005$$9785\pm.0007$$\mathbf{.0035\pm.0005}$$9784\pm.0007$$0037\pm.0007$$9785\pm.0007$$0055\pm.0017$$9785\pm.0005$
random forest .$1428\pm.0007$$9659\pm.0005$$0324\pm.0036$$9663\pm.0011$$\mathbf{.0052\pm.0018}$$9660\pm.0009$$0066\pm.0029$$9661\pm.0012$$0062\pm.0009$$9659\pm.0005$
1layer NN .$0164\pm.0004$$9760\pm.0005$$01$0023 .$9740\pm.0008$$0137\pm.0014$$9759\pm.0007$$0112\pm.0019$$9740\pm.0016$$\mathbf{.0062\pm.0026}$$9760\pm.0005$

### v2

@@ 4.2 Proposed Recalibration Algorithm Based on PAC-Bayes Bounds
From Corollary 2, we can see that as long as $\mathrm{KL}(\tilde{\rho}\|\tilde{\pi})$ is regularized and the ECE under the recalibration data is small, the recalibrated function could exhibit a small TCE. This leads us to propose a generalization-aware recalibration method by minimizing our bound.
However, minimizing this function is difficult since the ECE is neither smooth nor differentiable. To avoid this issue, we focus on the following relation:
$\mathrm{ECE}(\eta_{v}\circ f_{w},S_{\mathrm{re}})^{2}\leq\mathbb{E}_{(X,Y)\sim S_{\mathrm{re}}}\|\mathbf{e}_{Y}-\eta_{V}\circ f_{w}(X)\|_{2}^{2},$ (23)
in the multiclass classification, where $\mathbb{E}_{(X,Y)\sim S_{\mathrm{re}}}$ denotes the expectation by the empirical distribution of $S_{\mathrm{re}}$, see Appendix C.3 for its derivation. This upper bound is the Brier score (Gruber and Buettner, 2022), which is a continuous and differentiable calibration metric. Therefore, under the posterior $\tilde{\rho}(v)=\tilde{\rho}(v;\theta)$ parametrized by $\theta$, we can minimize the following PAC-Bayes-based objective by gradient-based optimization:
$\mathop{\mathbb{E}}\limits_{\tilde{\rho}(v;\theta)}\underbrace{\mathop{\mathbb{E}}\limits_{(X,Y)\sim S_{\mathrm{re}}}\|\mathbf{e}_{Y}\!-\eta_{V}\circ f_{w}(X)\|^{2}_{2}}_{\textrm{Brier score}}+\alpha\frac{\mathrm{KL}(\tilde{\rho}\|\tilde{\pi})}{n_{\mathrm{re}}}.$ (24)
@@
$\mathbb{E}_{\tilde{\rho}(v;\theta)}$ $\underbrace{\mathbb{E}_{(X,Y)\sim S_{\mathrm{re}}}l_{\mathrm{acc}}(Y,\eta_{V}\circ f_{w}(X))}_{\textrm{Expected classification loss}}$ (25)
$+\underbrace{\mathbb{E}_{(X,Y)\sim S_{\mathrm{re}}}\|\mathbf{e}_{Y}-\eta_{V}\circ f_{w}(X)\|_{2}^{2}}_{\textrm{Brier score}}+\alpha\frac{\mathrm{KL}(\tilde{\rho}\|\tilde{\pi})}{n_{\mathrm{re}}}.$ (26)
Since this objective function is derived from the PAC-Bayes bound for $l_{\textrm{acc}}$ (Theorem 4 in Appendix B) and Corollary 2 via a union bound, it remains within the generalization error bound (See Corollary 5 in Appendix C.3 for details).
We refer to the recalibration achieved by minimizing Eqs. (24) and (26) as PAC-Bayes recalibration (PBR) and summarize these procedures in Algorithm 1. As an example of a recalibration model in PBR, we use the Gaussian process (GP) (Rasmussen and Williams, 2005). In this case, $\tilde{\rho}(v;\theta)$ is set as the multivariate Gaussian distribution, $\mathcal{N}(v;\mu,\Sigma)$, where $\mu$ and $\Sigma$ are the mean and covariance matrix, respectively. Furthermore, we can construct a data-dependent GP prior $\tilde{\pi}$ using the $M$ outputs $\{f_{w}(x,i)\}_{i=1}^{M}$ ($x\in S_{\mathrm{re}}$) obtained during training as inducing points, which is expected to yield a small KL value. This is because $w$ is trained using $S_{\mathrm{tr}}$, and since $S_{\mathrm{tr}}$ and $S_{\mathrm{re}}$ are independent, this does not violate the independence of $S_{\mathrm{re}}$ from $\tilde{\pi}$ as assumed in Corollary 1.
A GP-based recalibration method was also proposed by Wenger et al. (2020), and, as far as we know, it is one of the methods that achieve state-of-the-art performance. Their recalibration was conducted via variational inference (VI), i.e., minimizing the following objective function derived from the evidence lower bound (ELBO):
$\mathbb{E}_{\tilde{\rho}(v;\theta)}\mathbb{E}_{(X,Y)\sim S_{\mathrm{re}}}l_{\mathrm{acc}}(Y,\eta_{V}\circ f_{w}(X))+\frac{\mathrm{KL}(\tilde{\rho}\|\tilde{\pi})}{n_{\mathrm{re}}},$
where $l_{\mathrm{acc}}$ is defined as the softmax loss. Eq. (26) corresponds to this ELBO objective when $\alpha=1$ and the Brier score is removed. This fact indicates that our PBR extends the GP-based recalibration to have flexible KL regularization and Brier score loss to control calibration performance on the basis of the PAC-Bayes notion. In this paper, we adopt in PBR the same computational strategy for GP-based recalibration as proposed by Wenger et al. (2020), in order to reduce the computational cost from $\mathcal{O}(N^{3})$ to $\mathcal{O}(N^{2}M)$. While this cost is not negligible, it remains justifiable in practice, as the size of $S_{\mathrm{re}}$ is significantly smaller than that of $S_{\mathrm{tr}}$. In Section 6.2, we compare PBR and GP-based recalibration using the same recalibration model settings.
@@
6:  (Option) Calculate the test ECE: $\mathrm{ECE}(\eta_{v}\circ f_{w},S_{\mathrm{te}})$ by setting $B=\lfloor n_{\mathrm{te}}^{1/3}\rfloor$.
7:  RETURN: $\eta_{v}\circ f_{w}$, ($\mathrm{ECE}(\eta_{v}\circ f_{w},S_{\mathrm{te}})$)
@@
Figure 1: KL vs ECE gap on MNIST (Multiclass; $1$-layer NN) and CIFAR-100 (Multiclass; ResNeXt-$29$ (Xie et al., 2017) and DenseNet-BC-190 (Huang et al., 2017)) dataset for each of various regularize parameters ($\alpha$).
@@
Table 1: Results of multiclass classification experiments on MNIST and CIFAR-100. The symbols ($\downarrow$) and ($\uparrow$) refer to lower and higher values indicating a higher performance, respectively. We set $n_{\mathrm{re}}=1000$.
Data Model Uncalibrate GP (Wenger et al., 2020) PBR (Ours; Eq. (24)) PBR (Ours; Eq. (26))) Temp. Scaling
@@
MNIST XGBoost .$0036\pm.0003$$9785\pm.0005$$0038\pm.0006$$9786\pm.0007$$\mathbf{.0035\pm.0004}$$9785\pm.0007$$0037\pm.0004$$9785\pm.0006$$0055\pm.0017$$9785\pm.0005$
random forest .$1428\pm.0007$$9659\pm.0005$$0313\pm.0056$$9663\pm.0011$$0065\pm.0017$$9639\pm.0018$$\mathbf{.0057\pm.0028}$$9656\pm.0017$$0062\pm.0009$$9659\pm.0005$
1layer NN .$0164\pm.0004$$9760\pm.0005$$01$0023 .$9740\pm.0008$$0137\pm.0014$$9751\pm.0009$$0112\pm.0019$$9740\pm.0016$$\mathbf{.0062\pm.0026}$$9760\pm.0005$
@@
5 Related work
Several studies have applied PAC-Bayes theory to classification metrics beyond accuracy. Ridgway et al. (2014) addressed AUC, Morvant et al. (2012) studied the confusion matrix in multiclass settings, and Sharma et al. (2024) examined conformal prediction as a tool for uncertainty quantification. In this work, we extend the PAC-Bayesian framework to the calibration metric.
Gupta and Ramdas (2021) and Kumar et al. (2019) examined the ECE in binary classification, focusing on uniform mass binning (UMB), which partitions the probability space into intervals of equal mass. They also analyzed a recalibration method based on UMB. Sun et al. (2023) further analyzed binning-based recalibration for binary classification under a similar Lipschitz assumption. Futami and Fujisawa (2024) investigated the generalization of the ECE using IT analysis under a comparable setting, and derived a total bias of order $\mathcal{O}(1/n^{1/3})$. In contrast, our work provides an analysis of the binning ECE in the multiclass setting—a direction that has not yet been explored. Furthermore, our bound is optimizable, which allows us to derive a novel recalibration algorithm for the ECE criterion that explicitly considers generalization. A natural direction for future work is to extend our analysis to settings that employ UMB.
The analysis of multiclass calibration remains limited compared to that of binary classification. For instance, Zhang et al. (2020) investigated the kernel-based $\mathrm{CE}_{K}$; however, their analysis does not address generalization or recalibration, which fundamentally distinguishes our approach. Moreover, while they noted that the binning-based $\mathrm{CE}_{K}$ suffers from the curse of dimensionality, no formal justification was provided. In contrast, our Theorem 3 formally validates this phenomenon. Gruber and Buettner (2022) discussed various calibration metrics and clarified that the TCE is a weaker calibration metric than $\mathrm{CE}_{K}$. Theorem 3 further demonstrates that the naive estimator of $\mathrm{CE}_{K}$ suffers more severely from the curse of dimensionality than the TCE.
Figure 2: $n_{\mathrm{re}}$ vs ECE on the test dataset on the experiments with CIFAR-100 dataset. PBR (ours) and PBR total (ours) correspond to the results obtained by optimizing Eqs. (24) and (26), respectively.
6 Experiments
In Section 6.1, we verify the correlation between our bounds in Corollary 1 and the generalization performance of recalibration. We then evaluate the effectiveness of our PBR described in Section 4.2. Due to page limitations, we primarily report results from multiclass classification experiments on the MNIST (LeCun et al., 1989) and CIFAR-100 (Krizhevsky, 2009) datasets. Additional experimental results are provided in Appendix E, and details of our experimental settings, including model specifications, are summarized in Appendix D. For ECE evaluation, we set $B=\lfloor n_{\mathrm{te}}^{1/3}\rfloor$ based on our findings.
6.1 Verification of Our Bounds
In this section, we empirically investigate the correlation between the KL divergence term in Eq. (19) and the ECE gap, which we define as the absolute difference between the ECEs of the test and training datasets. Our goal here is to assess whether the KL divergence in our bound can adequately explain the observed generalization performance. To investigate this, we first conducted recalibration experiments using PBR (Eq. (24)) with various KL constraints ($\alpha$). We then measured the KL values and the ECE gap after recalibration using $10$-fold cross-validation, setting $n_{\mathrm{re}}=1000$. As evaluation metrics, we adopted Pearson’s and Kendall’s rank-correlation coefficients, which are widely used to assess the consistency between generalization metrics and actual generalization performance (Jiang et al., 2020; Wang et al., 2021b; Kawaguchi et al., 2023).
A portion of the results is presented in Figure 1. These results confirm the positive correlation between the KL divergence and the ECE generalization gap. This confirms that our bounds are valid in explaining the generalization performance. It is also evident that the ECE gap is noisy. This is due to the slow convergence rate of the binning ECE estimator, which scales as $\mathcal{O}(1/n_{\mathrm{re}}^{1/3})$ [^1] We remark that we set $n_{\mathrm{te}}=n_{\mathrm{re}}$ when evaluating the ECE gap. even under the optimal $B$. With $n_{\mathrm{re}}=1000$, the estimation bias can contribute noise on the order of $\mathcal{O}(1/10)$, meaning that when models achieve an ECE below $0.1$, the correlation could be obscured by noise. A straightforward approach to reducing noise is increasing the size of the recalibration dataset. However, due to the $\mathcal{O}(1/n_{\mathrm{re}}^{1/3})$ convergence rate, this is impractical, highlighting the inherent difficulty of precise ECE evaluation.
6.2 Empirical Comparison of Recalibration Methods
In this section, we compare our proposed PBR and existing standard or state-of-the-art recalibration methods, temperature scaling (Guo et al., 2017) and GP calibration (Wenger et al., 2020). The optimizable parameter $\alpha$ in PBR is selected via grid search from $0\}$ , 0.01 , 0.1 , 0.25 , 0.5 , 0.75 , 0.9 , 1.0 }. Wenger et al. (2020) also reported results for recalibration methods originally developed for binary classification, such as Platt scaling (Platt, 1999), isotonic regression (Zadrozny and Elkan, 2002), Beta calibration (Kull et al., 2017), and Bayesian binning into quantiles (BBQ) (Naeini et al., 2015), by adapting these methods to the multiclass setting via one-vs-all conversion. To ensure a fair comparison, however, we included only methods that are directly compatible with multiclass classification as baselines. A comparison of these methods in the binary classification setting can be found in Appendix E.
We summarize the results in Table 1. The results show that our method consistently improves calibration performance over existing approaches, particularly outperforming the GP-based recalibration method. Additionally, we found that, contrary to the findings of Wenger et al. (2020), temperature scaling sometimes performs better than GP-based recalibration in terms of ECE when using the optimal $B$. One possible reason for this discrepancy is that Wenger et al. (2020) reported the ECE values with $B=100$. According to our analysis, this setting minimizes estimation bias when $n_{\mathrm{te}}=\mathcal{O}(100^{3})$. However, since the actual $n_{\mathrm{te}}$ used in their experiments is significantly smaller (see Section 4.1 of Wenger et al. (2020)), their reported results may have been affected by severe bias. This finding underscores the fundamental importance of selecting an appropriate number of bins based on a well-grounded theoretical framework for conducting reliable empirical evaluations.
These findings also indicate that the GP recalibration sometimes yield results that are not significantly different from the uncalibrated setting. In contrast, our PBR shows substantial improvements over both uncalibrated and GP-based methods. It is worth mentioning that our additional experiments in Appendix E confirm that GP-based methods, including PBR, do not consistently outperform standard recalibration techniques such as temperature scaling. These observations suggest that the use of GP for recalibration methods does not guarantee superior performance and emphasize the importance of carefully selecting a recalibration strategy that aligns with the characteristics of the dataset and model.
We also conducted additional experiments using $n_{\mathrm{re}}=\{10,50,100,250,500,1000,3000,5000\}$ to further examine how the performance of each recalibration method varies with different values of $n_{\mathrm{re}}$. The parameter $\alpha$ in PBR was selected following the same procedure as described earlier.
We show the results in Figure 2. From these results, we can see that our PBR achieves relatively good performance under various settings of $n_{\mathrm{re}}$, consistently improving the performance of GP-based recalibration methods. However, in experiments with AlexNet, GP-based recalibration including PBR fails except when $n_{\mathrm{re}}$ is small. Given AlexNet’s lower accuracy (see the Uncalibrated column in Table 1), the following explanation may be plausible. All of our recalibration methods construct the GP prior using the outputs of the pretrained model $f_{w}$ as inducing points. If $f_{w}$ fails to classify the training data accurately, the resulting GP prior will be misaligned with the true data distribution. Consequently, the recalibrated posterior is regularized toward this inappropriate prior, which can degrade overall performance. The temperature scaling appears to perform stable and effective recalibration as $n_{\mathrm{re}}$ increases, regardless of accuracy, and it even achieves a higher performance than PBR in some experimental settings; however, it is considered to cause overfitting when $n_{\mathrm{re}}$ is small because the temperature scaling lacks regularization terms such as the KL divergence used in PBR. This underscores the importance of carefully selecting recalibration methods based on the characteristics of the model and dataset as denoted previously. Summarizing these results, in practical applications—at least for multiclass classification—it would be advisable to try both temperature scaling and our PBR, verify their performance using validation data, and choose the method that contributes more to improving calibration performance.